OpenStudy (anonymous):

For position function s(t):=20-5t^2, if h is a small number, find the average rate of change of s(t) on the interval [2,2+h]. (Simplify fully)

OpenStudy (amistre64):

ug ..... i spose

OpenStudy (anonymous):

?

OpenStudy (amistre64):

s(2+h) -s(2) ----------- h

OpenStudy (amistre64):

its alot of work just to prove that derivatives work ....

OpenStudy (anonymous):

lol true.

OpenStudy (amistre64):

20-5(2+h)^2 -(20-5(2)^2) ------------------------ h 20-5(4 +4h +h^2) -20 +20 ------------------------ h 20-20-20h-5h^2 ---------------- h -20h-5h^2 ---------- = -20 -5h h

OpenStudy (anonymous):

I think I have to use the formula (f(a)-f(x))/(a-x)

OpenStudy (anonymous):

Can you explain to me how to use that formula?

OpenStudy (amistre64):

now for the derivative we get: -10t ; at t=2 we get -20 when h=0, we get -20

OpenStudy (amistre64):

its the same formula

OpenStudy (anonymous):

oh no derivatives yet

OpenStudy (anonymous):

can you explain to me how to use it? like what's a and what's x?

OpenStudy (anonymous):

on the interval

OpenStudy (amistre64):

yes, you are doing the proof for derivatives so that in the next few weeks your teacher will say, and this is how we do it the easy way...

OpenStudy (anonymous):

haha ok.

OpenStudy (amistre64):

f(a) and f(x) simply mean to use the formula with an a and an x instead of a t like this: f(a) - f(x) (20-5a^2) - (20-5x^2) ------------------- a-x

OpenStudy (amistre64):

20-5a^2 -20+5x^2 ----------------- a - x -5a^2+5x^2 ------------- a - x -5(a^2-x^2) ------------- a - x -5(a+x)(a-x) ------------- a - x -5(a+x)

OpenStudy (anonymous):

ok, what do you plug into f(a)?

OpenStudy (anonymous):

in [2,2+h]

OpenStudy (amistre64):

.............. a = 2 and x = h

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is a=2 from the 2+h part or from just the 2 part.

OpenStudy (amistre64):

it doesnt matter..... a =2 OR a=h, OR x=2 OR x=h.... -5(a+x) -5(2+h) is the same things as -5(h+2)

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

Like, I know the forumla, I just don't understand how to use it.

OpenStudy (amistre64):

then dont use it lol

OpenStudy (amistre64):

does it matter if f(a) = 2 and f(x) = (2+h)? no

OpenStudy (amistre64):

becasuse when h=0 it just disappears to -20

OpenStudy (amistre64):

\[\frac{y-y_0}{x-x_0}\] do you understand this formula?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

that is all it is; but they rename that parts ..... but it doesnt matter what they name them, they still do the same things

OpenStudy (anonymous):

so the bottom part is a-x... don't you put 2-h?

OpenStudy (anonymous):

why did you only put h?

OpenStudy (amistre64):

\[\frac{f(a)-f(x)}{a-x} \iff \frac{y-y_0}{x-x_0} \iff \frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

oh ok.

OpenStudy (amistre64):

because \((x+h)-(x)=h\)

OpenStudy (anonymous):

why?

OpenStudy (amistre64):

why what? why does x-x+h = 0+h = h?

OpenStudy (anonymous):

oh ok. just kidding.

OpenStudy (amistre64):

\[(x+h)-(x)\] \[x+h-x\] \[x-x+h\] \[(x-x)+h\] \[0+h\] \[h\]

OpenStudy (anonymous):

so the formula would be \[(f(2+h)-f(2))\div(h)\]

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so its ((20-5(2+h)^2)-(20-5(2)^2))/h

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so (20-5(2+h)^2)/h

OpenStudy (anonymous):

and finally -20-h?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

:) Thank you